OpenStudy (anonymous):

Sixty is 12% of what number? A.)5 B.)7.2 C.)500 D.)720 (please help me quickly!)

OpenStudy (anonymous):

60 = 0.12x ---> x = 500

OpenStudy (e.cociuba):

Let x = The number we are looking for. 12% of x = 60. 0.12x = 60 x = 60/0.12 x = 500 C.)

OpenStudy (anonymous):

60 = 12 %, what is 100 % = ? 60/12 = 12/12 % 5 = 1 %

OpenStudy (anonymous):

Oh okay so all I do is change 12 % into 0.12 witch equals 60 . so x=60/0.12 so x=500?

OpenStudy (e.cociuba):

Correct!

OpenStudy (anonymous):

Ah okay THANK YOU!

OpenStudy (e.cociuba):

Np :)